Erica flys her plane 200 mi against the wind in the same time it takes her to fly 300 mi with the wind. If the wind blows at 30mph what is the speed of the plane in still air?
Answer by jorel1380(3719)

200/x-30=300/x+30
200x+6000=300x-9000
15000=100x
150=x
The speed of the plane in still air is 150mph.